Bella Sandino made a $600 loan to her friend, Sadie, at 11%. Sadie paid back the loan with interest of $19.25. How long in days was the loan outstanding? (Use 360 days a year.)
Answer by Boreal(15235)

She paid back $619.25
619.25=600(1+r/n)=600(1+.11n/360); when n=360, the parentheses will equal 1.11
619.25/600=1+.11n/360
1.03208=1+.11/n
.03208=.11n/360
.11n=360*.03208=11.5488
n=104.99 days or 105 days ANSWER
This is between a 1/4 and 1/3 of a year, and 1 years' interest is $66, so the answer should be between $16.50 and $22.